# Scaling limits of critical FK-decorated random planar maps with $q=4$

### v1

@@ 2.2.2 Burger counts and scaling limits
In [SheffieldScott2016QGAI, Proposition 2.2], it is proved that for every $p\in[0,1]$, for the bi-infinite word $\ldots X(-1)X(0)X(1)\ldots$ with $(X(i))_{i\in\mathbb{Z}}$ i.i.d. copies of $X(1)$ with the law in (2.3), almost surely every letter has a match. For any portion $W$ of the word we define the burger discrepancy $\mathcal{D}(W)$ to be equal to $\mathsf{h}(W)-\#\mathsf{C}(W)-\#\mathsf{c}(W)+\#\mathsf{H}(W)$ plus one for every $\mathsf{F}$ symbol in $W$ whose match in the bi-infinite word is an $\mathsf{c}$, minus one for every $\mathsf{F}$ symbol in $W$ whose match in the bi-infinite word is an $\mathsf{h}$. Note that this match may not be in $W$, so the definition $\mathcal{D}(W)$ only makes sense for a portion $W$ of $\ldots X(-1)X(0)X(1)\ldots$.
@@
Sheffield [SheffieldScott2016QGAI] proved the following scaling limit result for the process defined by $(\mathcal{S}_{0},\mathcal{D}_{0})=(0,0)$ and
@@
Theorem 2.1 (Sheffield, [SheffieldScott2016QGAI]).
Let $p\in[0,1]$ and $\alpha=\max\{1-2p,0\}$. Let $B^{1}_{t}$ and $B^{2}_{t}$ be independent standard one-dimensional Brownian motion. Then
$\bigg(\frac{\mathcal{S}_{\lfloor nt\rfloor}}{\sqrt{n}},\frac{\mathcal{D}_{\lfloor nt\rfloor}}{\sqrt{n}}\bigg)_{t\in\mathbb{R}}\stackrel{{\scriptstyle\textnormal{d}}}{{\longrightarrow}}(B^{1}_{t},B^{2}_{\alpha t})$ (2.7)
@@
Observe that, when $p\geq 1/2$, the discrepancy goes to 0 at the scale $\sqrt{n}$. The case when $p>1/2$ is supposedly related to supercritical Liouville gravity, as it has been proved that the planar maps are tree-like [feng2023triviality]. Our main result, 1.1, establishes a non-trivial limit in the critical case $(q=4,p=1/2)$, when $\mathcal{D}_{n}$ is rescaled in the correct way.
@@ 2.2.3 Auxilliary walks: past and future
To analyse the word $X(\varphi(0))\cdots X(0)$, we introduce a pair of random walks $(h^{\leftarrow}_{n},c^{\leftarrow}_{n},n\geq 0)$ called the reduced walk (this terminology first appeared in [berestycki2017critical]), and sometimes referred to as the exploration into the past. [^3] These walks are denoted by $h$ and $c$ in [berestycki2017critical], but we use the left arrow to distinguish them from the exploration into the future, that we will define in the next paragraph and use in Section 6.
@@
  • if $Y(n)=\mathsf{H}$ (resp. $\mathsf{C}$) then $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1}+1,c^{\leftarrow}_{n-1})$ (resp. $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1},c^{\leftarrow}_{n-1}+1)$;
@@
  • if $Y(n)$ is an $\mathsf{F}$ excursion of type $\mathsf{c}$ (resp. type $\mathsf{h}$), then $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1}+|\overline{E}|,c^{\leftarrow}_{n-1})$ (resp. $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1},c^{\leftarrow}_{n-1}+|\overline{E}|)$) where $|\overline{E}|$ is the length of the reduced word. [^4] Note that if $E$ is an $\mathsf{F}$-excursion of type $\mathsf{c}$, the reduced word $\overline{E}$ can only consist of $\mathsf{H}$ orders. This explains why we consider these steps as increments for $h^{\leftarrow}$.
@@
The second exploration that will play a crucial role in the sequel will be looking into the future, i.e. to the right of 0. As far as we know, this exploration has not appeared in the literature before. Let $\tau_{\mathsf{F}}$ be the first time $k$ such that the reduced word $\overline{X(1,k)}$ contains an $\mathsf{F}$ symbol. In other words, $\tau_{\mathsf{F}}$ is the first time that we see an $\mathsf{F}$ symbol whose match lie at or to the left of position 0. From [SheffieldScott2016QGAI, Section 3.3], we know that $\tau_{\mathsf{F}}<\infty$ almost surely. Denote by $P_{\mathsf{F}}=X(1,\tau_{\mathsf{F}})$ the corresponding random word.
@@ 2.3.1 Definitions
Recall from Section 2.2 that the infinite FK planar map model is encoded by a bi-infinite word formed of symbols in $\{\mathsf{c},\mathsf{h},\mathsf{C},\mathsf{H},\mathsf{F}\}$, where each symbol $\mathsf{F}$ corresponds to a loop forming the interface between a primal and a dual connected component of the coloured (loop-decorated) triangulation associated to the map. We define a notion of typical loop by conditioning on the event $X(0)=\mathsf{F}$ and looking at the loop $\mathfrak{L}(0)$ corresponding to that $\mathsf{F}$ symbol. The loop is typical in the sense that a loop picked uniformly at random from a FK planar map of size $k$ converges in law to $\mathfrak{L}(0)$ as $k\to\infty$ (see [berestycki2017critical, Proposition 3.5]). There are also several equivalent definitions of a typical loop, as shown in [berestycki2017critical, Proposition 3.4]. For example, one could instead consider the first $\mathsf{F}$-excursion to the left of 0 (without any conditioning). The length of $\mathfrak{L}(0)$, denoted by $|\mathfrak{L}(0)|$, is defined as the number of triangles it crosses.
In [berestycki2017critical] and [GwynneEwain2019Slft], it is shown that for $q<4$ and $q=(2-4\cos^{2}(\frac{\pi}{4p_{0}}))^{2}$,
@@
In [GwynneEwain2019Slft], a slightly stronger result involving the regularity of the tails is obtained. In 1.3 we prove the analogous result for $q=4$ ($p_{0}=1/2$) and also identify the slowly varying correction exactly.
@@
When $X(0)=\mathsf{F}$, we also have the notion of envelope associated with $X(0)$ (the terminology “bubble” is also used interchangeably with envelope in [berestycki2017critical]). Recall that we call a word $e$ an $\mathsf{F}$-excursion if it is of the form $\mathsf{h}\cdots\mathsf{F}$ (type $\mathsf{h}$) or $\mathsf{c}\cdots\mathsf{F}$ (type $\mathsf{c}$), where the final $\mathsf{F}$ is matched to the first letter (Section 2.2). The associated envelope is then the (loop-decorated) submap of the infinite FK map encoded by this $\mathsf{F}$-excursion. We declare the root face of this submap to be one that is not crossed by a loop encoded by an $\mathsf{F}$ inside $e$. It also has simple boundary, and we denote by $|\partial\mathfrak{e}(0)|$ its boundary length. Conditioned on $X(0)=\mathsf{F}$, we write $\mathfrak{e}(0)$ for the envelope encoded by the $\mathsf{F}$-excursion $X(\varphi(0))\cdots X(0)$. Note that, conditional on $X(0)=\mathsf{F}$, both the typical loop and the typical cluster are included in the envelope $\mathfrak{e}(0)$. More precisely, since the match $\mathsf{c}\mathsf{F}$ (resp. $\mathsf{h}\mathsf{F}$) corresponds to a fictional edge in the encoding, $\mathsf{F}$-excursions of type $\mathsf{h}$ (resp. $\mathsf{c}$) correspond to envelopes with red outer boundary that surround a blue (resp. red) component, which is our typical cluster.
@@ 2.4 Fully packed loop-$O(n)$ model on triangulations
In this section we recall the fully packed loop-$O(n)$ model on planar triangulations and its connection to FK(q)(q)-decorated planar maps. The key feature of this model is that, via the gasket decomposition of [borot2012recursive], it satisfies a certain natural functional equation, which we will solve to compute the partition function at $n=2$.
@@
This model can be viewed as a symmetric case of the twofold loop model in [borot2012loop], where regions of the map separated by loops are assigned a colour red or blue depending only on the parity of the number of surrounding loops, and face/vertex weights may depend on the colour. In this picture, the boundary is monochromatic (since there is no loop visiting the root face) and the colouring procedure is completely determined by the colour of the boundary. We do not need the detailed coloured formulation here, only that when the face and vertex weights are chosen symmetrically this reduces to the fully packed $O(n)$ model above, and the partition functions corresponding to the two possible boundary colours coincide.
It is known [borot2012loop] that the Tutte map $T(\mathfrak{m},\mathfrak{p})$ of a self-dual FK(q)(q)-weighted map $(\mathfrak{m},\mathfrak{p})$ with $q=q(p)$ (as introduced in Section 2.1.1) has the law of a fully packed loop-$O(n)$ triangulation with weights
@@
$\Big(\frac{1-p}{16}\Big)^{{\frac{\#F(\mathfrak{t})+\ell-1}{2}}}\Big(\frac{2p}{1-p}\Big)^{\#\boldsymbol{\ell}+1}=nx_{c}^{\ell+1}x_{c}^{\#F(\mathfrak{t})-1}n^{\#\boldsymbol{\ell}}=nx_{c}^{\ell+1}Z(\mathfrak{t},\boldsymbol{\ell},n,x_{c}),$ (2.19)
@@
The gasket decomposition is a powerful approach towards the analysis of various loop models in [borot2012recursive], [borot2012loop]. We follow the exposition of [borot2012loop] but under the setting of our special fully packed model defined in (2.16) and (2.17).
@@
The contribution to the weight $Z({\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime}},x,n)$ in (2.16) of an internal face of degree $k$ in the gasket, whose associated ring of triangles has inner boundary length $k^{\prime}$ and $(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime})\in\mathbb{T}_{k^{\prime}}$ inside, is $nx^{k+k^{\prime}}Z(\mathfrak{t},\boldsymbol{\ell},x,n)$. We therefore define the weight
@@
It has been proved $W$ is analytic in $\mathbb{C}\setminus[a,b]$ where $[a,b]$ is a real interval with $b\geq|a|$ and $0\in[a,b]$. In this setting, [borot2012loop, Equation (3.22), (3.23)] simplify to
@@ 2.5 Proof outline
For context and comparison, we start by briefly recalling Sheffield’s argument [SheffieldScott2016QGAI] in the regime $0<q<4$. The proof is elementary in that it relies solely on martingale techniques. At a high level, the argument hinges upon two central ideas:
  • Tightness. The first key step is to show that the discrepancy $\mathcal{D}$ remains of the same order as the burger count $\mathcal{S}$ (denoted $\mathcal{C}$ in Sheffield’s notation). To this end, Sheffield introduces a clever lattice-path encoding of burger stacks, known as embedded stacks, which has the nice feature of being monotone [SheffieldScott2016QGAI, Section 3.4]. This monotonicity implies that $\mathcal{D}$ is at most of the same order as the burger count $\mathcal{S}$, leading to tightness in $n$ for the pair $(\mathcal{S}_{n}/\sqrt{n},\mathcal{D}_{n}/\sqrt{n})$.
@@
Instead, we use the planar map structure to obtain some preliminary information about the hamburger-cheeseburger walks. Our start point is to map the FK(4)(4) model to the loop-$O(2)$ model on triangulations, from which we derive both the exact expression and the asymptotic behaviour of the associated partition function $F_{\ell}$. This part, presented in Section 3, draws on ideas from analytic combinatorics (in particular the gasket decomposition introduced in [borot2012loop]) and has a different flavour from the rest of the article. It leads to the first rigorous derivation of the loop-$O(2)$ partition function expression and asymptotic (1.4), thereby confirming the physics prediction of Gaudin and Kostov [gaudin1989n].
Our strategy is then to approximate the discrepancy using Markovian walks, rather than dealing with the discrepancy directly. To do this, we construct two explorations – one into the past and one into the future – as defined in Section 2.2.3. These explorations admit a natural interpretation in terms of certain peeling explorations in the infinite-volume FK(4)(4) planar map. We relate in Section 4 some observables of the exploration into the past (namely, hitting times) to the partition function $F_{\ell}$, [^6] At a very high level, this part of the proof therefore bears some connections to the literature on the loop-$O(n)$ model on quadrangulations, where the partition function has been used to establish estimates on the peeling exploration [budd2018peeling, curien2019peeling]. which, in particular, implies the first two statements of 1.3. To the best of our knowledge, this connection appears to be novel. It goes through the notion of skeleton words that encode clusters through the Mullin–Bernardi–Sheffield bijection (Lemma 2.3). Using Tauberian theorems and regular variation techniques, it further enables us, in Section 5, to analyse the step distribution of the walk and to derive the scaling limit of the exploration into the past. We also emphasise that, because the regime $q=4$ is critical, both explorations will be proved to lie in the domain of attraction of a Cauchy-type process, which makes the analysis more subtle; in particular, we will make use of de Haan theory.
@@
Finally, we emphasise that, since we have access to an explicit expression for the partition function $F_{\ell}$, all our estimates are exact, capturing the integrable nature of the model. These techniques are further extended by Berestycki and Da Silva in a concurrent work to provide the exact critical exponents and partition functions of the FK(q)(q) and loop-$O(n)$ models in the regime $0<q<4$, $0<n<2$, which is the main focus of Sheffield’s work [SheffieldScott2016QGAI].

### v2

@@ 2.2.2 Burger counts and scaling limits
In [SHE16, Proposition 2.2], it is proved that for every $p\in[0,1]$, for the bi-infinite word $\ldots X(-1)X(0)X(1)\ldots$ with $(X(i))_{i\in\mathbb{Z}}$ i.i.d. copies of $X(1)$ with the law in (2.3), almost surely every letter has a match. For any portion $W$ of the word we define the burger discrepancy $\mathcal{D}(W)$ to be equal to $\mathsf{h}(W)-\#\mathsf{C}(W)-\#\mathsf{c}(W)+\#\mathsf{H}(W)$ plus one for every $\mathsf{F}$ symbol in $W$ whose match in the bi-infinite word is a $\mathsf{c}$, minus one for every $\mathsf{F}$ symbol in $W$ whose match in the bi-infinite word is an $\mathsf{h}$. Note that this match may not be in $W$, so the definition $\mathcal{D}(W)$ only makes sense for a portion $W$ of $\ldots X(-1)X(0)X(1)\ldots$.
@@
Sheffield [SHE16] proved the following scaling limit result for the process defined by $(\mathcal{S}_{0},\mathcal{D}_{0})=(0,0)$ and
@@
Theorem 2.1 (Sheffield, [SHE16]).
Let $p\in[0,1]$ and $\alpha=\max\{1-2p,0\}$. Let $B^{1}_{t}$ and $B^{2}_{t}$ be independent standard one-dimensional two-sided Brownian motions. Then
$\bigg(\frac{\mathcal{S}_{\lfloor nt\rfloor}}{\sqrt{n}},\frac{\mathcal{D}_{\lfloor nt\rfloor}}{\sqrt{n}}\bigg)_{t\in\mathbb{R}}\stackrel{{\scriptstyle\textnormal{d}}}{{\longrightarrow}}(B^{1}_{t},B^{2}_{\alpha t})_{t\in\mathbb{R}}$ (2.7)
@@
Observe that, when $p\geq 1/2$, the discrepancy goes to 0 at the scale $\sqrt{n}$. The case when $p>1/2$ is supposedly related to supercritical Liouville gravity, as it has been proved that the planar maps are tree-like [FEN25]. Our main result, 1.1, establishes a non-trivial limit in the critical case $(q=4,p=1/2)$, when $\mathcal{D}_{n}$ is rescaled in the correct way.
@@ 2.2.3 Auxilliary walks: past and future
To analyse the word $X(\varphi(0))\cdots X(0)$, we introduce a pair of random walks $(h^{\leftarrow}_{n},c^{\leftarrow}_{n},n\geq 0)$ called the reduced walk (this terminology first appeared in [BLR17]), and sometimes referred to as the exploration into the past. [^3] These walks are denoted by $h$ and $c$ in [BLR17], but we use the left arrow to distinguish them from the exploration into the future, that we will define in the next paragraph and use in Section 6.
@@
  • if $Y(n)=\mathsf{H}$ (resp. $\mathsf{C}$) then $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1}+1,c^{\leftarrow}_{n-1})$ (resp. $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1},c^{\leftarrow}_{n-1}+1)$);
@@
  • if $Y(n)$ is an $\mathsf{F}$ excursion $E$ of type $\mathsf{c}$ (resp. type $\mathsf{h}$), then $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1}+|\overline{E}|,c^{\leftarrow}_{n-1})$ (resp. $(h^{\leftarrow}_{n},c^{\leftarrow}_{n})=(h^{\leftarrow}_{n-1},c^{\leftarrow}_{n-1}+|\overline{E}|)$) where $|\overline{E}|$ is the length of the reduced word. [^4] Note that if $E$ is an $\mathsf{F}$-excursion of type $\mathsf{c}$, the reduced word $\overline{E}$ can only consist of $\mathsf{H}$ orders. This explains why we consider these steps as increments for $h^{\leftarrow}$.
@@
The second exploration that will play a crucial role in the sequel will be looking into the future, i.e. to the right of 0. As far as we know, this exploration has not appeared in the literature before. Let $\tau_{\mathsf{F}}$ be the first time $k$ such that the reduced word $\overline{X(1,k)}$ contains an $\mathsf{F}$ symbol. In other words, $\tau_{\mathsf{F}}$ is the first time that we see an $\mathsf{F}$ symbol whose match lie at or to the left of position 0. From [SHE16, Section 3.3], we know that $\tau_{\mathsf{F}}<\infty$ almost surely. Denote by $P_{\mathsf{F}}=X(1,\tau_{\mathsf{F}})$ the corresponding random word.
@@ 2.3.1 Definitions
Recall from Section 2.2 that the infinite FK planar map model is encoded by a bi-infinite word formed of symbols in $\{\mathsf{c},\mathsf{h},\mathsf{C},\mathsf{H},\mathsf{F}\}$, where each symbol $\mathsf{F}$ corresponds to a loop forming the interface between a primal and a dual connected component of the coloured (loop-decorated) triangulation associated to the map. We define a notion of typical loop by conditioning on the event $X(0)=\mathsf{F}$ and looking at the loop $\mathfrak{L}(0)$ corresponding to that $\mathsf{F}$ symbol. The loop is typical in the sense that a loop picked uniformly at random from a FK planar map of size $k$ converges in law to $\mathfrak{L}(0)$ as $k\to\infty$ (see [BLR17, Proposition 3.5]). There are also several equivalent definitions of a typical loop, as shown in [BLR17, Proposition 3.4]. For example, one could instead consider the first $\mathsf{F}$-excursion to the left of 0 (without any conditioning). The length of $\mathfrak{L}(0)$, denoted by $|\mathfrak{L}(0)|$, is defined as the number of triangles it crosses.
In [BLR17] and [GMS19], it is shown that for $q<4$ and $q=(2-4\cos^{2}(\frac{\pi}{4p_{0}}))^{2}$,
@@
In [GMS19], a slightly stronger result involving the regularity of the tails is obtained. In 1.3 we prove the analogous result for $q=4$ ($p_{0}=1/2$) and also identify the slowly varying correction exactly.
@@
When $X(0)=\mathsf{F}$, we also have the notion of envelope associated with $X(0)$ (the terminology “bubble” is also used interchangeably with envelope in [BLR17]). Recall that we call a word $e$ an $\mathsf{F}$-excursion if it is of the form $\mathsf{h}\cdots\mathsf{F}$ (type $\mathsf{h}$) or $\mathsf{c}\cdots\mathsf{F}$ (type $\mathsf{c}$), where the final $\mathsf{F}$ is matched to the first letter (Section 2.2). The associated envelope is then the (loop-decorated) submap of the infinite FK map encoded by this $\mathsf{F}$-excursion. We declare the root face of this submap to be one that is not crossed by a loop encoded by an $\mathsf{F}$ inside $e$. It also has simple boundary, and we denote by $|\partial\mathfrak{e}(0)|$ its boundary length. Conditioned on $X(0)=\mathsf{F}$, we write $\mathfrak{e}(0)$ for the envelope encoded by the $\mathsf{F}$-excursion $X(\varphi(0))\cdots X(0)$. Note that, conditional on $X(0)=\mathsf{F}$, both the typical loop and the typical cluster are included in the envelope $\mathfrak{e}(0)$. More precisely, since the match $\mathsf{c}\mathsf{F}$ (resp. $\mathsf{h}\mathsf{F}$) corresponds to a fictional edge in the encoding, $\mathsf{F}$-excursions of type $\mathsf{h}$ (resp. $\mathsf{c}$) correspond to envelopes with red outer boundary that surround a blue (resp. red) component, which is our typical cluster.
@@ 2.4 Fully packed loop-$O(n)$ model on triangulations
In this section we recall the fully packed loop-$O(n)$ model on planar triangulations and its connection to FK(q)(q)-decorated planar maps. The key feature of this model is that, via the gasket decomposition of [BBG11], it satisfies a certain natural functional equation, which we will solve to compute the partition function at $n=2$.
@@
This model can be viewed as a symmetric case of the twofold loop model in [BBG12a], where regions of the map separated by loops are assigned a colour red or blue depending only on the parity of the number of surrounding loops, and face/vertex weights may depend on the colour. In this picture, the boundary is monochromatic (since there is no loop visiting the root face) and the colouring procedure is completely determined by the colour of the boundary. We do not need the detailed coloured formulation here, only that when the face and vertex weights are chosen symmetrically this reduces to the fully packed $O(n)$ model above, and the partition functions corresponding to the two possible boundary colours coincide.
It is known [BBG12a] that the Tutte map $T(\mathfrak{m},\mathfrak{p})$ of a self-dual FK(q)(q)-weighted map $(\mathfrak{m},\mathfrak{p})$ with $q=q(p)$ (as introduced in Section 2.1.1) has the law of a fully packed loop-$O(n)$ triangulation with weights
@@
$\Big(\frac{1-p}{16}\Big)^{{\frac{\#F(\mathfrak{t})+\ell-1}{2}}}\Big(\frac{2p}{1-p}\Big)^{\#\boldsymbol{\ell}+1}=nx_{c}^{\ell+1}x_{c}^{\#F(\mathfrak{t})-1}n^{\#\boldsymbol{\ell}}=nx_{c}^{\ell+1}Z(\mathfrak{t},\boldsymbol{\ell},x_{c},n),$ (2.19)
@@
The gasket decomposition is a powerful approach towards the analysis of various loop models in [BBG11], [BBG12a]. We follow the exposition of [BBG12a] but under the setting of our special fully packed model defined in (2.16) and (2.17).
@@
The contribution to the weight $Z({\mathfrak{t},\boldsymbol{\ell}},x,n)$ in (2.16) of an internal face of degree $k$ in the gasket, whose associated ring of triangles has inner boundary length $k^{\prime}$ and $(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime})\in\mathbb{T}_{k^{\prime}}$ inside, is $nx^{k+k^{\prime}}Z(\mathfrak{t}^{\prime},\boldsymbol{\ell}^{\prime},x,n)$. We therefore define the weight
@@
It has been proved $W$ is analytic in $\mathbb{C}\setminus[a,b]$ where $[a,b]$ is a real interval with $b\geq|a|$ and $0\in[a,b]$. In this setting, [BBG12a, Equation (3.22), (3.23)] simplify to
@@ 2.5 Proof outline
For context and comparison, we start by briefly recalling Sheffield’s argument [SHE16] in the regime $0<q<4$. The proof is elementary in that it relies solely on martingale techniques. At a high level, the argument hinges upon two central ideas:
  • Tightness. The first key step is to show that the discrepancy $\mathcal{D}$ remains of the same order as the burger count $\mathcal{S}$ (denoted $\mathcal{C}$ in Sheffield’s notation). To this end, Sheffield introduces a clever lattice-path encoding of burger stacks, known as embedded stacks, which has the nice feature of being monotone [SHE16, Section 3.4]. This monotonicity implies that $\mathcal{D}$ is at most of the same order as the burger count $\mathcal{S}$, leading to tightness in $n$ for the pair $(\mathcal{S}_{n}/\sqrt{n},\mathcal{D}_{n}/\sqrt{n})$.
@@
Instead, we use the planar map structure to obtain some preliminary information about the hamburger-cheeseburger walks. Our start point is to map the FK(4)(4) model to the loop-$O(2)$ model on triangulations, from which we derive both the exact expression and the asymptotic behaviour of the associated partition function $F_{\ell}$. This part, presented in Section 3, draws on ideas from analytic combinatorics (in particular the gasket decomposition introduced in [BBG12a]) and has a different flavour from the rest of the article. It leads to the first rigorous derivation of the loop-$O(2)$ partition function expression and asymptotic (1.4), thereby confirming the physics prediction of Gaudin and Kostov [GK89].
Our strategy is then to approximate the discrepancy using Markovian walks, rather than dealing with the discrepancy directly. To do this, we construct two explorations – one into the past and one into the future – as defined in Section 2.2.3. These explorations admit a natural interpretation in terms of certain peeling explorations in the infinite-volume FK(4)(4) planar map. We relate in Section 4 some observables of the exploration into the past (namely, hitting times) to the partition function $F_{\ell}$, [^6] At a very high level, this part of the proof therefore bears some connections to the literature on the loop-$O(n)$ model on quadrangulations, where the partition function has been used to establish estimates on the peeling exploration [BUD18, CUR19]. which, in particular, implies the first two statements of 1.3. To the best of our knowledge, this connection appears to be novel. It goes through the notion of skeleton words that encode clusters through the Mullin–Bernardi–Sheffield bijection (Lemma 2.3). Using Tauberian theorems and regular variation techniques, it further enables us, in Section 5, to analyse the step distribution of the walk and to derive the scaling limit of the exploration into the past. We also emphasise that, because the regime $q=4$ is critical, both explorations will be proved to lie in the domain of attraction of a Cauchy-type process, which makes the analysis more subtle; in particular, we will make use of de Haan theory.
@@
Finally, we emphasise that, since we have access to an explicit expression for the partition function $F_{\ell}$, all our estimates are exact, capturing the integrable nature of the model. These techniques are further extended by Berestycki and Da Silva in the concurrent work [BD25] to provide the exact critical exponents and partition functions of the FK(q)(q) and loop-$O(n)$ models in the regime $0<q<4$, $0<n<2$, which is the main focus of Sheffield’s work [SHE16].